the perimeter of a rectangle is 20 inches. the length of the rectangle is 2inches longer than than the width.
write the equation to determine the width of the rectangle

Answer by oberobic(2304)

P = perimeter
L = length
W = width
P = 2(L+W)
.
P = 20
.
L = W+2
.
(W+2)*(W) = 20
.
W^2 + 2W = 20
.
Done.